# Byzantine Reliable Broadcast with Low Communication and Time Complexity

### v1

@@ 4.1. Overview
In this section, a variant of $\mathcal{A}_{bit}$, called $\mathcal{A}_{sig}$, is introduced with an optimal time complexity of $2$ while still guaranteeing that $\mathcal{L}(n,t)<3$. The challenge is that this time complexity cannot be reached when sending proposals to agree on a root hash first before sending fragments. However, if fragments are broadcast right away, a Byzantine sender may send fragments for different root hashes to different nodes. Naturally, honest nodes cannot simply broadcast any received fragments without an adverse effect on the communication complexity.
Algorithm $\mathcal{A}_{sig}$ addresses this challenge by using threshold signatures, i.e., we assume that each node has the keying material to threshold-sign messages and that every node can verify the validity of signatures of all nodes. At least $12t+$ 1 threshold signatures are required to create a valid signature. The proposals of $\mathcal{A}_{bit}$ are replaced by messages containing a root hash and a threshold signature or a full signature. The crucial advantage is that a (full) signature serves as proof that the sender has received threshold signatures from $12t+$ 1 nodes, ensuring that the execution can make progress whenever an honest nodes meets the conditions to deliver $m$.
@@ 4.2. Description
     if valid_multi_sig($h,\sigma$) then
        $h^{*}\coloneqq h$
     else if (($|H(v_{k})|<2$ or $h\in H(v_{k})$) and valid_sig($h,v_{k},\sigma$then
@@
Algorithm 3 Algorithm $\mathcal{A}_{sig}$: Triggered actions at node $v_{i}$. Initially, $F=R=H=S=\{\}$ and $h^{*}=\bot$.
Algorithm $\mathcal{A}_{sig}$ uses signed proposal messages of the form $proposal(h,\sigma)$, where $\sigma$ is either the threshold signature of $h$ of a particular node or it is a signature of $h$ derived from at least $12t+$ 1 threshold signatures. Note that it is a signature of $h$ and the identifier in practice in order to tie the signature to a particular execution. The algorithm still uses the hash maps $F$, $H$, and $R$. Threshold signatures are collected in the hash map $S:\mathcal{H}\times V\rightarrow\mathcal{S}$, where $\mathcal{S}$ denotes the set of all possible signatures. We further define that $S(h)\coloneqq\cup_{v\in V}S(h,v)$. In this section, $h_{max}$ is defined as $h_{max}\coloneqq\arg\max_{h\in\mathbf{H}}|S(h)|$. In addition to $done$, each node further stores $h^{*}$, the root hash of the message that will be delivered if set (initially, $h^{*}=\bot$).
The steps of algorithm $\mathcal{A}_{sig}$ for triggered actions are shown in Algorithm 3. When reliable_broadcast is invoked, exactly the same steps as in Algorithm 1 are executed. A message $fragment(h,j,f_{j},\pi_{j})$ is handled the same way as in $\mathcal{A}_{bit}$ with the exception that $h$ is signed and both $proposal(h,\sigma_{i})$ and $fragment(h,f_{i},i,\pi_{i})$ are broadcast if it is the first (valid) fragment received from $sender$. When receiving $proposal(h,\sigma)$ from some node $v_{k}$, the action depends on $\sigma$. If it is a valid signature, $h^{*}$ is set to $h$. Otherwise, if it is a valid threshold signature and $H(v_{k})$ does not contain two hashes different from $h$, then $h$ is added to $H(v_{k})$ and $S(h,v_{k})$ is set to $\sigma$.
The state-based actions at node $v_{i}$ are shown in Algorithm 4. If at least $12t+$ 1 threshold signatures have been collected for some root hash $h_{max}$, then $h^{*}$ is set to $h_{max}$. Node $v_{i}$ broadcasts its fragment $f_{i}$ (with the Merkle proof $\pi_{i}$) if $h^{*}$ is set, $f_{i}$ is locally available, and it has not been broadcast before. Lastly, if $h^{*}\neq\bot$ and $|F(h^{*})|\geq 2t+1$, $v_{i}$ computes the signature and broadcasts the corresponding proposal before executing the same delivery steps as in Algorithm 2.
@@ 4.3. Analysis
According to Algorithm 3, only the first message from $sender$ is threshold-signed, i.e., every honest node provides a threshold signature for at most one hash. However, setting $h^{*}$ to some hash $h$ requires $12t+$ 1 threshold signatures of $h$. If we assume that there are two different hashes for which $12t+$ 1 threshold signatures have been collected, there must be at least one honest node that threshold-signed two different hashes, which contradicts the rule that honest nodes only threshold-sign at most one hash. ∎
@@
     $\sigma\coloneqq$ compute_signature($S(h^{*})$)
     broadcast $proposal(h^{*},\sigma)$
@@
Algorithm 4 Algorithm $\mathcal{A}_{sig}$: State-based actions at node $v_{i}$. Initially, $F=R=S=\{\}$, $h^{*}=\bot$, and $done=false$. Let $h_{max}\coloneqq\arg\max_{h\in\mathcal{H}}|S(h)|$.
@@
An honest node $v_{i}$ broadcasts its fragment $f_{i}$ for some root hash $h$ when receiving $f_{i}$ from $sender$. Apart from this transmission, according to Algorithm 3, $v_{i}$ only sends fragments associated with $h^{*}$. Since $h^{*}$ never changes once set, the claim follows. ∎
@@
Algorithm $\mathcal{A}_{sig}$ and $\mathcal{A}_{bit}$ share the property that messages for two different hashes are accepted from any node. According to Algorithm 3, an honest node broadcasts a signature and fragment message for the same hash $h$ when receiving its fragment from $sender$. Algorithm 4 states that any further transmission of a signature or fragment message must be for hash $h^{*}\neq\bot$. Since $h^{*}$ never changes when set, honest nodes send messages for at most $2$ hashes. ∎
Algorithm $\mathcal{A}_{sig}$ depends on the same mechanism to ensure totality as $\mathcal{A}_{bit}$ and therefore the following variant of Lemma 3.3 holds.
Lemma 4.4 ().
@@
The same argument as for algorithm $\mathcal{A}_{bit}$ applies, which proves that every honest node gets its fragment eventually after an honest node $v$ has delivered a message $m$ for some hash $h$, which it accepts and stores due to Lemma 4.3. Since $v$ must have computed and broadcast a valid signature $\sigma$, it eventually holds at all honest nodes that $h^{*}=h$. According to Lemma 4.1, honest nodes all set $h^{*}$ to the same value. Consequently, every honest node will eventually broadcast its fragment, Thus, every honest node will eventually have at least $12t+$ 1 fragments of $m$. ∎
As in §3, the first main result is that algorithm $\mathcal{A}_{sig}$ is a correct reliable broadcast algorithm.
Theorem 4.5 ().
@@
The four conditions of reliable broadcast are again proved in sequence.
Validity: If the sender $v$ of a message $m$ with root hash $h$ is honest, it sends $f_{j}$ to $v_{j}$ for all $v_{j}\in V$. Every honest node $v_{j}$ broadcasts its fragment and threshold signature and, consequently, receives at least $12t+$ 1 signatures and fragments eventually. Subsequently, every honest node computes the signature $\sigma$, sets $h^{*}\coloneqq h$, and proceeds to deliver $m$.
@@
Totality: If an honest node $v$ delivers a message $m$ with root hash $h$, it eventually holds at all honest nodes that $h^{*}=h$ and $|F(h^{*})|\geq 2t+1$ due to Lemma 4.4. Since $v$ managed to verify the correctness of the fragments and the root hash, it holds again that the verification must succeed at other nodes as well. Every honest node delivers $m$ after the successful verification. ∎
@@
Let $S$ denote the set of honest nodes that send fragments for more than one root hash. Algorithm 3 dictates that an honest node only sends another fragment when $h^{*}\neq\bot$. If there are $t^{\prime}\leq t$ Byzantine nodes in the execution, at least $2t+1-t^{\prime}$ honest nodes must receive their fragments for hash $h^{*}$ in order to obtain a signature $\sigma$ for this hash. Since none of these nodes sends fragments for other root hashes, we get that $|S|\leq 3t+1-(2t+1-t^{\prime})-t^{\prime}=t$. ∎
As the following theorem shows, algorithm $\mathcal{A}_{sig}$ achieves an overhead factor that tends to $5/2$ for large messages.
@@
It holds for algorithm $\mathcal{A}_{sig}$ that $\mathcal{L}(n,t)=\frac{5}{2}$.
@@
The dissemination of fragments by the sender requires at most $(n-1)|f|$ bits, where $|f|$ again denotes the size of the largest fragment message sent by an honest node. Due to Lemma 4.6, at most $t$ honest nodes broadcast fragments twice, and the other other honest nodes broadcast at most one fragment. Additionally, each node sends $O(\kappa n)$ bits for the signed proposal messages and $f_{j}$ to all $v_{j}\in V\setminus R(h)$, where $|V\setminus R(h)|\leq t<n/3$ as before. Hence it follows that $\mathcal{C}(n,t,|m|)$ is at most
@@
where the first inequality holds because $(5t+1)<\frac{5}{3}(3t+1)=\frac{5}{3}n$. Thus, $\mathcal{L}(n,t)=\frac{5}{2}$. ∎
It is worth noting that the communication complexity is only worse compared to $\mathcal{A}_{bit}$ if the sender deliberately transmits conflicting fragments, i.e., $\mathcal{L}(n,t)=2$ still holds otherwise. The main advantage of $\mathcal{A}_{sig}$ over $\mathcal{A}_{bit}$ is its superior time complexity.
@@
If the sender is honest, all honest nodes receive their fragments within 1 time unit and broadcast their threshold signatures and fragments. After another time unit, all honest nodes must have received at least $12t+$ 1 threshold signatures and fragments for the same root hash. As a result, each honest node computes the signature and sets $h^{*}$, which triggers the delivery of $m$. ∎
@@
According to Lemma 4.6, there are at most $t$ honest nodes that send their fragments for $2$ root hashes, i.e., there are at least $t+1$ honest nodes that only send their fragment once. Thus, an honest node may store $2t\cdot 2+(t+1)\cdot 1\leq\frac{5}{3}n$ fragments of size $\frac{3}{2}\ell_{max}/n$ for a total of $\frac{5}{3}\ell_{max}$.
Algorithm $\mathcal{A}_{sig}$ also utilizes $R$, containing at most $2$ node identifiers of size $O(\kappa)$ per node, and $S$, storing at most $2$ signatures of size $O(\kappa)$ per node. Thus, the space complexity of these data structures is upper bounded by $O(n\kappa)$. ∎
@@ 5. Practical Considerations
Partially synchronous communication. Byzantine behavior and unpredictable message latencies are often exceptional situations rather than the norm in practice. While both algorithm $\mathcal{A}_{sig}$ and $\mathcal{A}_{bit}$ are tailored to the asynchronous communication model, they can easily be adapted to the partially synchronous model where periods of synchrony are assumed. Given an upper bound $d$ on the message delay that holds for some periods of time, both algorithms can be modified to improve the communication complexity during these times in the absence of faults. Specifically, the additional constraint can be introduced that at least $\delta$ time must have passed since the first fragment was received before delivering $m$. The following theorem states the effect of this modification on the communication complexity of both algorithms for an algorithm-specific $\delta$.
Theorem 5.1 ().
If communication is synchronous from the start of the execution for at least $\delta=3d$ ($\delta=2d$) time, then $\mathcal{L}(n,t)=\frac{3}{2}$ for the adapted version of $\mathcal{A}_{bit}$ ($\mathcal{A}_{sig}$) in the absence of faults.
@@
As there are no faults by assumption, Theorem 3.6 and Theorem 4.8 imply that all nodes receive $n$ fragments after at most $3d$ ($2d$) time in case of algorithm $\mathcal{A}_{bit}$ ($\mathcal{A}_{sig}$). Thus, $R(h)=V$ for the root hash $h$ of $m$, which entails that the nodes do not disseminate any additional fragments before delivering, resulting in a communication complexity of $(n-1)|f|+n(n-1)|f|+O(\kappa n)<n^{2}|f|+O(\kappa n)\leq\frac{3}{2}|m|n+O(\kappa n^{2}\log(n))$ and thus $\mathcal{L}(n,t)=\frac{3}{2}$. ∎
Naturally, this modification does not improve any theoretical (worst-case) bound because it still holds that $\mathcal{L}(n,t)=2$ even for $t$ crash failures but it may prove to be beneficial for real-world applications where failures are rare nonetheless.
Simplified structure. The second optimization only concerns $\mathcal{A}_{sig}$. Instead of handling fragments and proposals in separate messages, signatures can be appended to the fragment message. As a result, nodes only send and process messages of a single type. Fragments and proposals are kept separate in algorithm $\mathcal{A}_{sig}$ to retain as much similarity to $\mathcal{A}_{bit}$ as possible in order to emphasize the key differences and simplify the presentation. This modification requires the addition of a simple rule: a node appends its threshold signature as long as $h^{*}=\bot$ and the signature otherwise, which can be computed if $h^{*}\neq\bot$ according to Algorithm 4. The signature in the message is still processed as shown in Algorithm 3.
Algorithm 4 dictates that the signature is broadcast before executing the message delivery. This step can be omitted entirely when adding the rule above without violating any of the reliable broadcast conditions. Since the signature is broadcast to ensure totality, we briefly argue why this condition continues to hold. Let $v$ be an honest node that has delivered message $m$ with root hash $h$. If all honest nodes broadcast fragments for root hash $h$ at some point, then all honest nodes will eventually deliver $m$ even if they never receive a signature. On the other hand, if there is at least one honest node $v^{\prime}$ that never sends a fragment for the right hash, then $v^{\prime}\notin R(h)$ at node $v$, and thus $v$ must send $v^{\prime}$’s fragment and the signature to $v^{\prime}$. Upon receipt of this message, $v^{\prime}$ will broadcast its fragment and the signature, ensuring that all honest nodes eventually receive the signature, i.e., the signature is broadcast anyway in this case.
6. Lower Bound
@@
In the following, we consider the case $r=3$, which permits a simple strategy to derive a lower bound due to the fact that a node failing to receive expected messages in the second round can only inform other nodes about the failure in the third round. Requiring a fourth round to send the missing bits violates the requirement that all honest nodes deliver the message by round $3$ in case of a non-faulty sender. Therefore, the strategy is simply to mark those nodes as faulty that send many bits in rounds $2$ and $3$.
Given this strategy, it is easy to see that $|S_{j}|>t$ and $\bar{b}_{j}>0$ must hold for any $v\in V\setminus R$ as otherwise $v_{j}$ may not receive $m$ by the end of round $3$. The following lemma provides a lower bound on $\bar{b}_{j}$ for the case $|S_{j}|>t$.
Lemma 6.1 ().
@@
bits by the end of round $3$, which implies that $v_{j}$ does not receive the (whole) message. ∎
@@
[…]

### v2

@@ 4.1. Overview
In this section, a variant of $\mathcal{A}_{bit}$, called $\mathcal{A}_{sig}$, is introduced with an optimal time complexity of $2$ while still guaranteeing that $\mathcal{L}(n)<3$. The challenge is that this time complexity cannot be reached when sending proposals to agree on a root hash first before sending fragments. However, if fragments are broadcast right away, a Byzantine sender may send fragments for different root hashes to different nodes. Naturally, honest nodes cannot simply broadcast any received fragments without an adverse effect on the communication complexity.
Algorithm $\mathcal{A}_{sig}$ addresses this challenge by using threshold signatures, i.e., we assume that each node has the keying material to threshold-sign messages and that every node can verify the validity of signature shares and signatures of all nodes. At least $12t+$ 1 signature shares are required to create a valid signature. The proposals of $\mathcal{A}_{bit}$ are replaced by messages containing a root hash and a signature share or a signature. The crucial advantage is that a signature serves as proof that the sender has received signature shares from $12t+$ 1 nodes, ensuring that the execution can make progress whenever an honest nodes meets the conditions to deliver $m$.
@@ 4.2. Description
Algorithm $\mathcal{A}_{sig}$ uses signed proposal messages of the form $proposal(h,\sigma)$, where $\sigma$ is either a signature share of $h$ of a particular node or a signature of $h$ derived from at least $12t+$ 1 signature shares. Note that, in practice, it is a signature of $h$ and the identifier in order to tie the signature to a particular execution. The algorithm still uses the hash maps $F$, $H$, and $R$. Signature shares are collected in the hash map $S:\mathcal{H}\times V\rightarrow\mathcal{S}$, where $\mathcal{S}$ denotes the set of all possible signature shares. We further define that $S(h)\coloneqq\cup_{v\in V}S(h,v)$. In this section, $h_{max}$ is defined as $h_{max}\coloneqq\arg\max_{h\in\mathbf{H}}|S(h)|$. In addition to $done$, each node further stores $h^{*}$, the root hash of the message that will be delivered if set, and the corresponding signature $\sigma^{*}$ (initially, $h^{*}=\sigma^{*}=\bot$).
@@
     if valid_signature($h,\sigma$) then
        $\sigma^{*}\coloneqq\sigma$, $h^{*}\coloneqq h$
     else if (($|H(v_{k})|<2$ or $h\in H(v_{k})$) and valid_share($h,v_{k},\sigma$then
@@
Algorithm 3 $\mathcal{A}_{sig}$: Triggered actions at node $v_{i}$. Initially, $F=R=H=S=\{\}$, $h^{*}=\sigma^{*}=\bot$.
The steps of algorithm $\mathcal{A}_{sig}$ for triggered actions are shown in Algorithm 3. When reliable_broadcast is invoked, exactly the same steps as in Algorithm 1 are executed. A fragment message is handled the same way as in $\mathcal{A}_{bit}$ apart from two differences: $h$ is signed and both $proposal(h,\sigma_{i})$ and $fragment(h,f_{i},i,\pi_{i})$ are broadcast if it is the first (valid) fragment received from the sender. When receiving $proposal(h,\sigma)$ from some node $v_{k}$, the action depends on $\sigma$. If it is a valid signature, $\sigma^{*}$ and $h^{*}$ are set to $\sigma$ and $h$, respectively. Otherwise, if it is a valid signature share and $H(v_{k})$ does not contain two hashes different from $h$, then $h$ is added to $H(v_{k})$ and $S(h,v_{k})$ is set to $\sigma$.
The state-based actions are shown in Algorithm 4. If at least $12t+$ 1 signature shares have been collected for some root hash $h_{max}$, then $\sigma^{*}$ is set to the computed signature and $h^{*}$ is set to $h_{max}$. Node $v_{i}$ broadcasts $f_{i}$ (including $\pi_{i}$) if $h^{*}$ is set, $f_{i}$ is locally available, and it has not been broadcast before. Lastly, if $h^{*}\neq\bot$ and $|F(h^{*})|\geq 2t+1$, $v_{i}$ broadcasts $proposal(h^{*},\sigma^{*})$ before executing the same delivery steps as in Algorithm 2.
@@ 4.3. Analysis
According to Algorithm 3, only the first message from the $sender$ is threshold-signed, i.e., every honest node provides a signature share for at most one hash. However, setting $h^{*}$ to some hash $h$ requires $12t+$ 1 signature shares of $h$. If we assume that there are two different hashes for which $12t+$ 1 signature shares have been collected, there must be at least one honest node that threshold-signed two different hashes, which contradicts the rule that honest nodes only threshold-sign at most one hash. ∎
@@
     broadcast $proposal(h^{*},\sigma^{*})$
@@
Algorithm 4 $\mathcal{A}_{sig}$: State-based actions at node $v_{i}$. Initially, $F=R=S=\{\}$, $h^{*}=\sigma^{*}=\bot$, and $done=false$. Let $h_{max}\coloneqq\arg\max_{h\in\mathcal{H}}|S(h)|$.
@@
An honest node $v_{i}$ broadcasts its fragment $f_{i}$ for some root hash $h$ when receiving $f_{i}$ from the $sender$. Apart from this transmission, according to Algorithm 3, $v_{i}$ only sends fragments associated with $h^{*}$. Since $h^{*}$ never changes once set, the claim follows. ∎
@@
Algorithm $\mathcal{A}_{sig}$ and $\mathcal{A}_{bit}$ share the property that messages for two different hashes are accepted from any node. According to Algorithm 3, an honest node broadcasts a signature and fragment message for the same hash $h$ when receiving its fragment from the $sender$. Algorithm 4 states that any further transmission of a signature or fragment message must be for hash $h^{*}\neq\bot$. Since $h^{*}$ never changes when set, the claim follows. ∎
The following variant of Lemma 3.3 holds for algorithm $\mathcal{A}_{sig}$.
Lemma 4.4.
@@
The same argument as for algorithm $\mathcal{A}_{bit}$ applies, which proves that every honest node gets its fragment eventually after an honest node $v$ has delivered a message $m$ for some hash $h$, which it accepts due to Lemma 4.3. Since $v$ must have computed and broadcast a valid signature $\sigma$, it eventually holds at all honest nodes that $h^{*}=h$ (and $\sigma^{*}=\sigma$). According to Lemma 4.1, honest nodes all set $h^{*}$ to the same value. Consequently, every honest node will eventually broadcast its fragment and thus also receive at least $12t+$ 1 fragments of $m$. ∎
As in §3, the first main result is that $\mathcal{A}_{sig}$ is a correct reliable broadcast algorithm.
Theorem 4.5.
@@
The four conditions of reliable broadcast are again proved separately.
Validity: If the sender $v$ of a message $m$ with root hash $h$ is honest, it sends $f_{j}$ to $v_{j}$ for all $v_{j}\in V$. Every honest node $v_{j}$ broadcasts its fragment and signature share and, consequently, receives at least $12t+$ 1 signature shares and fragments eventually. It follows that every honest node eventually computes $\sigma^{*}$, sets $h^{*}\coloneqq h$, and proceeds to deliver $m$.
@@
Totality: If an honest node $v$ delivers a message $m$ with root hash $h$, it eventually holds at all honest nodes that $h^{*}=h$ and $|F(h^{*})|\geq 2t+1$ due to Lemma 4.4. Since $v$ managed to verify the correctness of the fragments and the root hash, it holds that the verification must succeed at other nodes as well, which entails that every honest node delivers $m$. ∎
@@
Let $S$ denote the set of honest nodes that send fragments for more than one root hash. Algorithm 3 dictates that honest nodes only send another fragment when $h^{*}\neq\bot$. If there are $t^{\prime}\leq t$ Byzantine nodes in the execution, at least $2t+1-t^{\prime}$ honest nodes must receive their fragments for hash $h^{*}$ in order to obtain a signature $\sigma^{*}$ for this hash. Since none of these nodes sends fragments for other root hashes, we get that $|S|\leq 3t+1-(2t+1-t^{\prime})-t^{\prime}=t$. ∎
As the following theorem shows, $\mathcal{A}_{sig}$ achieves an overhead factor of $5/2$ for $|m|\rightarrow\infty$.
@@
It holds for algorithm $\mathcal{A}_{sig}$ that $\mathcal{L}(n)=\frac{5}{2}$.
@@
The dissemination of fragments by the sender requires at most $(n-1)|f|$ bits, where $|f|$ again denotes the size of the largest fragment message sent by an honest node. Due to Lemma 4.6, at most $t$ honest nodes broadcast fragments twice, and the other honest nodes broadcast at most one fragment. Additionally, each node sends $O(\kappa n)$ bits for the signed proposal messages and $f_{j}$ to all $v_{j}\in V\setminus R(h)$, where $|V\setminus R(h)|\leq t<n/3$ as before. Hence it follows that $\mathcal{C}(n,|m|)$ is at most
@@
where the first inequality holds because $(5t+1)<\frac{5}{3}(3t+1)=\frac{5}{3}n$. Thus, $\mathcal{L}(n)=\frac{5}{2}$. ∎
It is worth noting that the communication complexity is only worse compared to $\mathcal{A}_{bit}$ if the sender deliberately transmits conflicting fragments, i.e., $\mathcal{L}(n)=2$ still holds otherwise. The main advantage of $\mathcal{A}_{sig}$ over $\mathcal{A}_{bit}$ is its superior time complexity.
@@
If the sender is honest, all honest nodes receive their fragments within $1$ time unit and broadcast their signature shares and fragments. After $2$ time units, all honest nodes must have received at least $12t+$ 1 signature shares and fragments for the same root hash. Thus, all honest nodes compute the signature $\sigma^{*}$ and set $h^{*}$, triggering the delivery of $m$. ∎
@@
According to Lemma 4.6, there are at most $t$ honest nodes that send their fragments for $2$ root hashes, i.e., there are at least $t+1$ honest nodes that only send their fragment once. Thus, an honest node may store $2t\cdot 2+(t+1)\cdot 1\leq\frac{5}{3}n$ fragments of size $\frac{3}{2}\ell_{max}/n$ for a total of $\frac{5}{2}\ell_{max}$. Algorithm $\mathcal{A}_{sig}$ also utilizes $R$, containing at most $2$ node identifiers of size $O(\kappa)$ per node, and $S$, storing at most $2$ signature shares of size $O(\kappa)$ per node. Thus, the space complexity of these data structures is upper bounded by $O(n\kappa)$. ∎
@@ 5. Practical Considerations
Partially synchronous communication. Byzantine behavior and unpredictable message latencies are often exceptional situations in practice. While $\mathcal{A}_{sig}$ and $\mathcal{A}_{bit}$ are both tailored to the asynchronous communication model, they can easily be adapted to the partially synchronous model where periods of synchrony are assumed. Given an upper bound $d$ on the message delay that holds for some periods of time, both algorithms can be modified to improve the communication complexity during these times in the absence of faults. Specifically, the additional constraint can be introduced that at least $\delta$ time must have passed since the first fragment was received before delivering $m$. The following theorem states the effect of this modification on the communication complexity for an algorithm-specific $\delta$.
Theorem 5.1.
If communication is synchronous from the start of the execution for $\delta=3d$ ($\delta=2d$) time and there are no faults, then $\mathcal{L}(n)=\frac{3}{2}$ for the adapted version of $\mathcal{A}_{bit}$ ($\mathcal{A}_{sig}$).
@@
As there are no faults by assumption, Theorem 3.6 and Theorem 4.8 imply that all nodes receive $n$ fragments after at most $3d$ ($2d$) time for $\mathcal{A}_{bit}$ ($\mathcal{A}_{sig}$). Thus, $R(h)=V$ for the root hash $h$ of $m$, which entails that the nodes do not disseminate any additional fragments before delivering, resulting in a communication complexity of $(n-1)|f|+n(n-1)|f|+O(\kappa n)<n^{2}|f|+O(\kappa n)\leq\frac{3}{2}|m|n+O(\kappa n^{2}\log(n))$ and thus $\mathcal{L}(n)=\frac{3}{2}$. ∎
Naturally, this modification does not improve any worst-case bounds because $\mathcal{L}(n)=2$ still holds even for $t$ crash failures but it may be beneficial in practice nonetheless.
Simplified structure. The second optimization only concerns $\mathcal{A}_{sig}$. Instead of handling fragments and proposals separately, signature shares and signatures can be appended to the fragment message. As a result, nodes only send and process messages of a single type. Fragments and proposals are kept separate in algorithm $\mathcal{A}_{sig}$ to retain as much similarity to $\mathcal{A}_{bit}$ as possible in order to emphasize the key differences and simplify the presentation. This modification requires the addition of a simple rule: a node appends its signature share as long as $h^{*}=\bot$ and the signature $\sigma^{*}$ otherwise. Note that $\sigma^{*}\neq\bot$ always holds when $h^{*}\neq\bot$. Signature shares and signatures are still processed as shown in Algorithm 3.
According to Algorithm 4, the signature is broadcast before executing the message delivery. This step can be omitted entirely when adding the rule above without violating any of the reliable broadcast conditions. Since the signature is broadcast to ensure totality, we briefly argue why this condition continues to hold. Let $v$ be an honest node that has delivered message $m$ with root hash $h$. If all honest nodes broadcast fragments for root hash $h$ at some point, then all honest nodes will eventually deliver $m$ even if they never receive a signature. On the other hand, if there is at least one honest node $v^{\prime}$ that never sends a fragment for the right hash, then $v^{\prime}\notin R(h)$ at node $v$, and thus $v$ must send $v^{\prime}$’s fragment and the signature to $v^{\prime}$. Upon receipt of this message, $v^{\prime}$ will broadcast its fragment and the signature, ensuring that all honest nodes eventually receive the signature, i.e., the signature is broadcast anyway in this case.
@@
In the following, we consider the case $r=3$, which permits a simple strategy to derive a lower bound because a node failing to receive expected messages in the second round can only inform other nodes about the failure in the third round. Requiring a fourth round to send the missing bits violates the requirement that all honest nodes deliver the message by round $3$ in case of a non-faulty sender. Therefore, the strategy is simply to mark those nodes as faulty that send many bits in rounds $2$ and $3$. Given this strategy, it is easy to see that $|S_{j}|>t$ and $\bar{b}_{j}>0$ must hold for any $v_{j}\in V\setminus R$ as otherwise $v_{j}$ may not receive $m$ by the end of round $3$. The following lemma provides a lower bound on $\bar{b}_{j}$ for the case $|S_{j}|>t$.
Lemma 5.2.
@@
bits by the end of round $3$, which implies that $v_{j}$ does not receive the whole message. ∎
@@
$\geq\sum_{v_{j}\in V\setminus R}\frac{n}{n-t}(|m|-b_{j})=(n-|R|)\frac{n}{n-t}|m|-\sum_{v_{j}\in V\setminus R}\frac{n}{n-t}b_{j}$
(1)${{\geq}}(n-|R|)\frac{n}{n-t}|m|-o(|m|n)\stackrel{{\scriptstyle n\rightarrow\infty}}{{=}}n\frac{n}{n-n/3}|m|=\frac{3}{2}|m|n.\qed$